sin 0 sec 0 = cos 0 csc 0 4) %3D
7) 6 sin2x - 5 sin x - 1 = 0
Answer #1

gino feco= CofO. Csco fino L Cose Sine rano= Cot O 6sin a- 5 Sina-1=0 6sin x-6 sinxt fina-=0 6 fina (sina-i]41 ina-1) = 0 (si